# Towards Understanding Link Predictor Generalizability Under Distribution Shifts

### v2

@@ Appendix D Additional Analysis Details
Split Train/Test Train/Valid Valid/Test Original 83.50 ± 7.33 83.40 ± 7.33 86.91 ± 6.58 CN (0,1,2) 87.73 ± 6.12 87.42 ± 6.17 91.78 ± 4.01 (0,2,4) 87.06 ± 6.08 86.72 ± 6.14 91.01 ± 4.29 (0,3,5) 86.45 ± 6.12 86.19 ± 6.12 90.42 ± 4.41 (2,1,0) 85.53 ± 6.74 85.22 ± 6.79 90.45 ± 4.49 (4,2,0) 85.97 ± 7.09 85.55 ± 7.15 90.71 ± 4.43 (5,3,0) 86.08 ± 7.29 85.62 ± 7.34 90.74 ± 4.53 SP ($\infty$,6,4) 84.11 ± 6.94 82.12 ± 7.34 85.60 ± 6.75 ($\infty$,4,3) 84.35 ± 6.93 82.66 ± 7.45 85.97 ± 6.75 (4,6$\infty$) 81.10 ± 7.74 82.15 ± 7.79 81.86 ± 7.48 (3,4$\infty$) 81.99 ± 7.84 82.81 ± 7.98 82.76 ± 7.72 PA (0,50,100) 83.92 ± 7.22 82.59 ± 7.63 85.50 ± 7.14 (0,100,200) 87.48 ± 6.16 87.55 ± 6.10 92.52 ± 3.70 (0,150,250) 87.61 ± 6.10 87.81 ± 6.17 92.45 ± 3.84 (100,50,0) 82.28 ± 7.96 83.22 ± 8.00 83.17 ± 7.85 (200,100,0) 84.75 ± 6.65 85.32 ± 6.64 89.94 ± 4.52 (250,150,0) 85.08 ± 6.65 85.51 ± 6.64 90.62 ± 4.46
Appendix E Time to Split Tested Dataset Samples
Split ogbl-collab ogbl-ppa ogbl-ddi CN (0, 1, 2) 7.48 s 177.89 s 13.45 s (0, 2, 4) 7.49 s 177.09 s 14.21 s (0, 3, 5) 7.63 s 178.23 s 13.64 s (2, 1, 0) 7.66 s 184.98 s 13.31 s (4, 2, 0) 7.63 s 186.49 s 13.71 s (5, 3, 0) 7.86 s 185.95 s 13.42 s SP ($\infty$, 6, 4) 53.12 s 2748.64 s ($\infty$, 4, 3) 52.24 s 2705.91 s (4, 6, $\infty$) 53.93 s 2715.34 s (3, 4, $\infty$) 53.7 s 2751.56 s PA (0, 50, 100) 19.25 s 406.5 s 15.41 s (0, 100, 200) 19.05 s 408.04 s 14.55 s (0, 150, 250) 19.35 s 407.81 s 13.96 s (100, 50, 0) 19.65 s 425.3 s 14.42 s (200, 100, 0) 19.34 s 409.93 s 14.59 s (250, 150, 0) 19.37 s 403.55 s 14.05 s
Appendix F Size of Dataset Samples
In this section we detail the number of training, validation, and test edges for all of the newly created splits detailed in Section 3. There are in Tables 6 for ogbl-collab, ogbl-ppa, and ogbl-ddi respectively.
Table 6. Number of samples in the ogbl-collab, ogbl-ppa, and ogbl-ddi datasets for the forward and backward heuristic splits, with shared split columns across datasets.
Split ogbl-collab ogbl-ppa ogbl-ddi Train Valid Test Train Valid Test Train Valid Test CN (0, 1, 2) 57638 6920 4326 2325936 87880 67176 3938 1184 212 (0, 2, 4) 237928 20045 14143 3000000 95679 83198 5636 992 525 (0, 3, 5) 493790 31676 21555 3000000 98081 88778 6580 942 809 (2, 1, 0) 1697336 23669 9048 3000000 96765 92798 10000 966 1808 (4, 2, 0) 1193456 24097 11551 3000000 93210 85448 10000 1668 2296 (5, 3, 0) 985820 25261 11760 3000000 92403 81887 10000 1462 2832 SP ($\infty$, 6, 4) 46880 1026 2759 17464 149 20 - - - ($\infty$, 4, 3) 52872 1238 3457 134728 4196 1180 - - - (4, 6, $\infty$) 1882392 5222 2626 3000000 90511 458 - - - (3, 4, $\infty$) 1877626 4384 7828 3000000 97121 74068 - - - PA (0, 50, 100) 210465 46626 9492 95671 95671 45251 5132 1572 419 (0, 100, 200) 329383 62868 25527 98562 98562 63178 8246 3208 1422 (0, 150, 250) 409729 75980 39429 99352 99352 72382 11718 5058 3694 (100, 50, 0) 1882392 64729 41381 3000000 90623 44593 10000 2678 1900 (200, 100, 0) 1877626 64202 30983 3000000 89671 34321 10000 6188 3632 (250, 150, 0) 457372 65323 30999 3000000 91995 35088 10000 6300 5970
Appendix G How effectively does LPShift induce distribution shift?
The following section will explore the capability of LPShift to induce a measurably-significant distribution shift in the structure of the ogbl-collab dataset. We apply the 2-sample Kolmgorov-Smirnov (KS) test (Hodges Jr, 1958) to compare if training, validation, and testing distributions can be sampled from one another, both before and after applying LPShift. As a controlled baseline to test LPShift, we generate a Holme-Kim (HK) graph (Holme and Kim, 2002) with a 40% chance to close a triangle, allowing the HK graph to contain numerous Common Neighbors without becoming fully-connected. HK graph generation parameters are included in Appendix D.
@@ Appendix G How effectively does LPShift induce distribution shift?
Appendix H Dataset Results
In this section, we include all of the results for each experiment conducted on the LPShift dataset splits.
@@ Appendix H Dataset Results
Table 10. ogbl-collab results on the forward and backward splits. Results are reported for DropEdge and TC (BUDDY) and for GCN+BUDDY and RA+BUDDY (Filter+Rank).
Split BUDDY Filter+Rank DropEdge TC GCN+BUDDY RA+BUDDY CN (0, 1, 2) 15.54 ± 0.98 11.27 ± 2.03 8.50 ± 1.10 3.94 ± 0.51 (0, 2, 4) 16.16 ± 0.17 12.39 ± 1.43 12.85 ± 0.83 6.61 ± 0.30 (0, 3, 5) 16.34 ± 0.17 14.99 ± 1.69 15.35 ± 0.96 7.16 ± 0.09 (2, 1, 0) 2.61 ± 0.09 5.28 ± 0.05 5.46 ± 0.11 4.32 ± 0.19 (4, 2, 0) 2.88 ± 0.11 5.23 ± 0.05 5.36 ± 0.12 4.45 ± 0.12 (5, 3, 0) 4.99 ± 0.08 6.03 ± 0.10 5.96 ± 0.06 4.93 ± 0.11 SP ($\infty$, 6, 4) 12.31 ± 0.51 7.25 ± 0.81 7.38 ± 0.82 3.84 ± 0.59 ($\infty$, 4, 3) 17.11 ± 1.02 9.33 ± 1.66 9.60 ± 0.39 7.63 ± 0.25 (4, 6, $\infty$) 5.34 ± 0.43 6.88 ± 0.30 6.86 ± 1.13 6.86 ± 1.13 (3, 4, $\infty$) 2.93 ± 0.18 6.24 ± 0.13 6.47 ± 0.24 6.47 ± 0.24 PA (0, 50, 100) 21.35 ± 0.36 18.82 ± 1.35 15.92 ± 1.01 15.92 ± 1.01 (0, 100, 200) 13.84 ± 0.64 12.13 ± 1.04 9.47 ± 0.31 9.47 ± 0.31 (0, 150, 250) 12.85 ± 0.78 10.63 ± 0.55 9.60 ± 0.41 9.60 ± 0.41 (100, 50, 0) 26.09 ± 0.62 15.84 ± 1.13 14.34 ± 1.05 14.34 ± 1.05 (200, 100, 0) 15.68 ± 0.85 9.15 ± 0.39 8.35 ± 0.34 8.35 ± 0.34 (250, 150, 0) 13.13 ± 0.94 6.70 ± 0.20 5.50 ± 0.33 5.50 ± 0.33
Table 11. ogbl-ppa results on the forward and backward splits. Results are reported for DropEdge and TC (BUDDY) and for GCN+BUDDY and RA+BUDDY (Filter+Rank).
Split BUDDY Filter+Rank DropEdge TC GCN+BUDDY RA+BUDDY CN (0, 1, 2) $7.83\pm 0.27$ $5.27\pm 0.34$ $4.48\pm 0.33$ $4.04\pm 0.26$ (0, 2, 4) $3.83\pm 0.25$ $2.91\pm 0.06$ $3.79\pm 0.28$ $3.42\pm 0.20$ (0, 3, 5) $3.06\pm 0.06$ $2.67\pm 0.13$ $3.16\pm 0.10$ $2.95\pm 0.13$ (2, 1, 0) $1.61\pm 0.04$ $3.44\pm 0.08$ $3.19\pm 0.08$ $2.58\pm 0.09$ (4, 2, 0) $2.47\pm 0.07$ $3.45\pm 0.10$ $3.25\pm 0.09$ $3.04\pm 0.05$ (5, 3, 0) $2.56\pm 0.08$ $3.55\pm 0.13$ $3.36\pm 0.13$ $3.10\pm 0.15$ SP ($\infty$, 6, 4) $3.86\pm 0.39$ $4.00\pm 0.29$ $4.00\pm 0.20$ $3.89\pm 0.23$ ($\infty$, 4, 3) $5.87\pm 0.16$ $4.82\pm 0.39$ $5.53\pm 0.92$ $5.53\pm 0.94$ (4, 6, $\infty$) $3.86\pm 0.39$ $13.2\pm 0.45$ $14.41\pm 0.67$ $13.63\pm 0.97$ (3, 4, $\infty$) $5.87\pm 0.16$ $2.89\pm 0.09$ $2.93\pm 0.15$ $2.51\pm 0.14$ PA (0, 5k, 10k) $3.93\pm 0.98$ $3.62\pm 0.21$ $3.66\pm 0.38$ $3.66\pm 0.38$ (0, 10k, 20k) $6.38\pm 3.48$ $3.13\pm 0.12$ $3.19\pm 0.22$ $3.19\pm 0.22$ (0, 15k, 25k) $2.49\pm 0.01$ $3.33\pm 1.38$ $2.88\pm 0.06$ $2.88\pm 0.06$ (10k, 5k, 0) $3.13\pm 0.10$ $1.78\pm 0.16$ $2.05\pm 0.05$ $2.05\pm 0.04$ (20k, 10k, 0) $2.56\pm 0.19$ $1.50\pm 0.0008$ $1.67\pm 0.03$ $1.65\pm 0.05$ (25k, 15k, 0) $2.40\pm 0.03$ $1.53\pm 0.04$ $1.66\pm 0.02$ $1.64\pm 0.03$
Table 12. ogbl-ddi results on the forward and backward splits. Results are reported for DropEdge and TC (BUDDY) and for GCN+BUDDY and RA+BUDDY (Filter+Rank).
Split BUDDY Filter+Rank DropEdge TC GCN+BUDDY RA+BUDDY CN (0, 1, 2) 1.65 ± 0.17 1.1± 0.01 0.87 ± 0.21 0.87 ± 0.21 (0, 2, 4) 2.75 ± 0.16 1.01 ± 0.01 1.45 ± 0.35 1.45 ± 0.35 (0, 3, 5) 2.15 ± 0.08 0.94 ± 0.01 1.65 ± 0.22 1.65 ± 0.22 (2, 1, 0) 0.78 ± 0.06 1.04 ± 0.02 1.20 ± 0.14 1.20 ± 0.14 (4, 2, 0) 1.09 ± 0.22 0.9 ± 0.02 2.44 ± 0.27 2.44 ± 0.27 (5, 3, 0) 1.40 ± 0.35 0.98 ± 0.03 3.19 ± 0.36 3.19 ± 0.36 PA (0, 5k, 10k) 1.81 ± 0.35 1.04 ± 0.01 1.26 ± 0.28 1.26 ± 0.28 (0, 10k, 20k) 1.74 ± 0.28 1.06 ± 0.02 1.43 ± 0.1 1.43 ± 0.11 (0, 15k, 25k) 1.92 ± 0.08 1.24 ± 0.02 1.04 ± 0.23 1.04 ± 0.23 (10k, 5k, 0) 3.83 ± 3.22 1.0 ± 0.03 9.16 ± 9.27 9.16 ± 9.27 (20k, 10k, 0) 2.71 ± 1.60 1.66 ± 0.01 9.84 ± 1.97 9.84 ± 1.97 (25k, 15k, 0) 11.07 ± 1.44 1.95 ± 0.02 9.50 ± 2.18 9.50 ± 2.18
@@
Table 13. MRR Results for traditional generalization methods applied to GCN on the Forward splits for ogbl-collab. Generalization methods that significantly improve performance are bolded.
@@
Table 14. MRR Results for traditional generalization methods applied to GCN on the Forward splits for ogbl-ppa. Generalization methods that significantly improve performance are bolded.
@@
Table 15. MRR Results for traditional generalization methods applied to GCN on the Forward splits for ogbl-ddi. Generalization methods that significantly improve performance are bolded.
@@ Appendix I Earth Mover’s Distance (EMD) Results
Appendix J Additional Training Details
Please consult the project README for building the project, loading data, and re-creating results. Tuned model hyperparameters are further detailed within their respective run scripts.
Table 17. Fixed Model Hyperparameters by model. Note: Hidden channels were fixed across all datasets.
GCN BUDDY NCNC NeoGNN LPFormer SEAL Hidden Channels 128 256 256 256 128 256
  • All models and link predictors are each fixed to 3 layers in their neural architecture.
  • We apply the ’NCNC2’ variant of NCNC with an added depth argument of 2 (Wang et al., 2023a) for all CN and SP splits of ogbl-collab. Otherwise, we set depth equal to 1, so as to reduce runtime.
  • NeoGNN use 1-hop neighborhoods on all tested datasets.
[…]
Appendix K Dataset Licenses
@@
Appendix M Impact Statement

### v3

@@ Appendix E Is LPShift effective at inducing distribution shift?
Appendix F Dataset Results
This section includes Tables 9 to 11, displaying baseline results for initial LPShift experiments.
@@ Appendix F Dataset Results
Appendix G Generalization Results
This section includes Tables 12 to 14, displaying results for the graph-specific generalization methods conducted on the LPShift dataset splits.
Table 12. ogbl-collab (MRR) results under LPShift after applying graph-specific generalization methods.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) DropEdge 15.54 ± 0.98 16.16 ± 0.17 16.34 ± 0.17 2.61 ± 0.09 2.88 ± 0.11 4.99 ± 0.08 TC 11.27 ± 2.03 12.39 ± 1.43 14.99 ± 1.69 5.28 ± 0.05 5.23 ± 0.05 6.03 ± 0.10 GCN+BUDDY 8.50 ± 1.10 12.85 ± 0.83 15.35 ± 0.96 5.46 ± 0.11 5.36 ± 0.12 5.96 ± 0.06 RA+BUDDY 3.94 ± 0.51 6.61 ± 0.30 7.16 ± 0.09 4.32 ± 0.19 4.45 ± 0.12 4.93 ± 0.11 SP ($\infty$,6,4) ($\infty$,4,3) (4,6$\infty$) (3,4$\infty$) DropEdge 12.31 ± 0.51 17.11 ± 1.02 5.34 ± 0.43 2.93 ± 0.18 TC 7.25 ± 0.81 9.33 ± 1.66 6.88 ± 0.30 6.24 ± 0.13 GCN+BUDDY 7.38 ± 0.82 9.60 ± 0.39 6.86 ± 1.13 6.47 ± 0.24 RA+BUDDY 3.84 ± 0.59 7.63 ± 0.25 6.86 ± 1.13 6.47 ± 0.24 PA (0,50,100) (0,100,200) (0,150,250) (100,50,0) (200,100,0) (250,150,0) DropEdge 21.35 ± 0.36 13.84 ± 0.64 12.85 ± 0.78 26.09 ± 0.62 15.68 ± 0.85 13.13 ± 0.94 TC 18.82 ± 1.35 12.13 ± 1.04 10.63 ± 0.55 15.84 ± 1.13 9.15 ± 0.39 6.70 ± 0.20 GCN+BUDDY 15.92 ± 1.01 9.47 ± 0.31 9.60 ± 0.41 14.34 ± 1.05 8.35 ± 0.34 5.50 ± 0.33 RA+BUDDY 15.92 ± 1.01 9.47 ± 0.31 9.60 ± 0.41 14.34 ± 1.05 8.35 ± 0.34 5.50 ± 0.33
Table 13. ogbl-ppa (MRR) results under LPShift after applying graph-specific generalization methods.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) DropEdge 7.83 ± 0.27 3.83 ± 0.25 3.06 ± 0.06 1.61 ± 0.04 2.47 ± 0.07 2.56 ± 0.08 TC 5.27 ± 0.34 2.91 ± 0.06 2.67 ± 0.13 3.44 ± 0.08 3.45 ± 0.10 3.55 ± 0.13 GCN+BUDDY 4.48 ± 0.33 3.79 ± 0.28 3.16 ± 0.10 3.19 ± 0.08 3.25 ± 0.09 3.36 ± 0.13 RA+BUDDY 4.04 ± 0.26 3.42 ± 0.20 2.95 ± 0.13 2.58 ± 0.09 3.04 ± 0.05 3.10 ± 0.15 SP ($\infty$,6,4) ($\infty$,4,3) (4,6$\infty$) (3,4$\infty$) DropEdge 3.86 ± 0.39 5.87 ± 0.16 3.86 ± 0.39 5.87 ± 0.16 TC 4.00 ± 0.29 4.82 ± 0.39 13.2 ± 0.45 2.89 ± 0.09 GCN+BUDDY 4.00 ± 0.20 5.53 ± 0.92 14.41 ± 0.67 2.93 ± 0.15 RA+BUDDY 3.89 ± 0.23 5.53 ± 0.94 13.63 ± 0.97 2.51 ± 0.14 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) DropEdge 3.93 ± 0.98 6.38 ± 3.48 2.49 ± 0.01 3.13 ± 0.10 2.56 ± 0.19 2.40 ± 0.03 TC 3.62 ± 0.21 3.13 ± 0.12 3.33 ± 1.38 1.78 ± 0.16 1.50 ± 0.0008 1.53 ± 0.04 GCN+BUDDY 3.66 ± 0.38 3.19 ± 0.22 2.88 ± 0.06 2.05 ± 0.05 1.67 ± 0.03 1.66 ± 0.02 RA+BUDDY 3.66 ± 0.38 3.19 ± 0.22 2.88 ± 0.06 2.05 ± 0.04 1.65 ± 0.05 1.64 ± 0.03
Table 14. ogbl-ddi (MRR) results under LPShift after applying graph-specific generalization methods.
CN (0,1,2) (0,2,4) (0,3,5) (2,1,0) (4,2,0) (5,3,0) DropEdge 1.65 ± 0.17 2.75 ± 0.16 2.15 ± 0.08 0.78 ± 0.06 1.09 ± 0.22 1.40 ± 0.35 TC 1.10 ± 0.01 1.01 ± 0.01 0.94 ± 0.01 1.04 ± 0.02 0.90 ± 0.02 0.98 ± 0.03 GCN+BUDDY 0.87 ± 0.21 1.45 ± 0.35 1.65 ± 0.22 1.20 ± 0.14 2.44 ± 0.27 3.19 ± 0.36 RA+BUDDY 0.87 ± 0.21 1.45 ± 0.35 1.65 ± 0.22 1.20 ± 0.14 2.44 ± 0.27 3.19 ± 0.36 PA (0,5k,10k) (0,10k,20k) (0,15k,25k) (10k,5k,0) (20k,10k,0) (25k,15k,0) DropEdge 1.81 ± 0.35 1.74 ± 0.28 1.92 ± 0.08 3.83 ± 3.22 2.71 ± 1.60 11.07 ± 1.44 TC 1.04 ± 0.01 1.06 ± 0.02 1.24 ± 0.02 1.00 ± 0.03 1.66 ± 0.01 1.95 ± 0.02 GCN+BUDDY 1.26 ± 0.28 1.43 ± 0.10 1.04 ± 0.23 9.16 ± 9.27 9.84 ± 1.97 9.50 ± 2.18 RA+BUDDY 1.26 ± 0.28 1.43 ± 0.11 1.04 ± 0.23 9.16 ± 9.27 9.84 ± 1.97 9.50 ± 2.18
Appendix H Traditional OOD Generalization Method Results
This section includes table detailing the performance of domain-agnostic/traditional OOD generalization methods on all of LPShift’s tested dataset splits.
@@ Appendix H Traditional OOD Generalization Method Results
Table 15. MRR Results for traditional OOD generalization methods on LPShift’s ogbl-collab.
@@
Table 16. MRR Results for traditional OOD generalization methods on LPShift’s ogbl-ppa.
@@
Table 17. MRR Results for traditional OOD generalization methods on LPShift’s ogbl-ddi.
@@ Appendix I Earth Mover’s Distance (EMD) Results
This section contains the EMD calculations for all dataset splits and a visualization for the EMD calculations on the ”Forward” ogbl-ddi dataset split.
@@
Appendix J Dataset Licenses
@@
Appendix L Impact Statement